OpenStudy (anonymous):

What is the value of csc (2pi/3)?

OpenStudy (jdoe0001):

what's the \(\bf sin\left(\frac{2\pi}{3}\right)\) ?

OpenStudy (jdoe0001):

csc = 1/sin so whatever \(\bf sin\left(\frac{2\pi}{3}\right)\) is, just flip that upside down

OpenStudy (anonymous):

Oh okay thanks!